# Geometric QCD II: The Confining Twistor String and Meson Spectrum

### v1

@@ 13.5 Balance of quark forces and boundary condition
This result confirms that the linear confinement arises directly from the kinematics of the Hodge-dual minimal surface in Minkowski space, without invoking independent string vibrations or ad-hoc constants.
14 Twistor Holography and the Rigid Twistor String
14.1 Twistor Holography (Gauge Holography)
14.2 Analytic Twistor String vs. Topological Twistor String
The general issues of the discretization of our theory and its direct numerical simulation (DNS) are discussed in D, where we present formulas and algorithms for the analytic extension of holomorphic twistor fields from a discrete set of boundary values. We attempted the direct numerical simulation of this path integral with oscillating weight using the Complex Langevin equation (CLE) on E. The results are encouraging but not convincing; perhaps some additional numerical algorithms are needed to overcome the instabilities of the CLE and obtain reliable results.
15 Numerical Simulation of Confining String spectrum
The analytic twistor-string formulation of the previous sections reduces the meson spectroscopy problem to the analytic structure of the quark two-point amplitude $\Pi(E)$ (after analytic continuation to Minkowski energy $E$). Since the corresponding boundary-twistor path integral is not localized, the evaluation of the spectrum becomes primarily computational. In this section, we summarize the numerical strategy and present preliminary proof-of-concept results. Full implementation details, discretization formulas, drift terms, and numerical diagnostics are given in E.
Spectral observables and mass extraction
The physical quark two-point amplitude $\Pi(E)$ is a meromorphic function of the energy, characterized by physical poles at the bound-state meson masses. However, directly searching for vertical asymptotes (poles) in a stochastic Monte Carlo simulation is a numerical trap: statistical variance near a singularity diverges uncontrollably, producing massive fat-tail excursions and destroying the Complex Langevin evolution.
To completely bypass this fatal instability, we apply an analytic transformation that converts the dangerous meromorphic observable into a perfectly bounded, smooth entire function. Rather than evaluating $\Pi(E)$ directly, we consider the logarithmic derivative
$R(E)=\frac{\Pi^{\prime}(E)}{\Pi(E)},$ (15.1)
and reconstruct the spectrum from the zero-crossings of its exact inverse, $1/R(E)$.
This inversion maps vertical asymptotes into well-behaved roots. Near a physical mass pole, $\Pi(E)\sim(E-m_{n})^{-1}$, the inverse resolvent becomes $1/R(E)\sim m_{n}-E$. The singularity is mathematically flattened into a smooth zero-crossing with a strictly negative slope. Conversely, near an artifact zero of the vacuum polarization ($\Pi(E)\sim E-E_{0}$), one obtains $1/R(E)\sim E-E_{0}$, yielding a zero-crossing with a strictly positive slope.
Because $1/R(E)$ is an entire function with no singularities anywhere in the complex plane (as detailed in Appendix D.7), the Complex Langevin process can sample it safely. Therefore, the $\bar{q}q$ masses are unambiguously identified as the zero-crossings of $\text{Re}[1/R(E)]$ possessing a negative slope. Furthermore, because the true physical signal is bounded, we know a priori that any massive spikes in our simulation data are purely statistical fat-tail excursions, rigorously justifying aggressive outlier filtering (such as median pooling) without fear of clipping physical resonances.
Reweighted Complex Langevin strategy (summary)
The effective twistor-string functional integral has an oscillatory weight; at present, Complex Langevin evolution (CLE) provides a practical route to evaluating it. A key technical point is that pole-finding is numerically unstable in a stochastic computation. We convert the problem into a robust zero-finding problem by a reweighted formulation, summarized here and derived in Appendix D.
Let $S_{\rm eff}$ denote the effective action for the discretized boundary-twistor variables at fixed $E$, and define the energy-derivative observable
$O\;\equiv\;\frac{\partial S_{\rm eff}}{\partial E}\,.$ (15.2)
Then one has the exact identity
$\frac{1}{R(E)}\;=\;\frac{\Pi(E)}{\Pi^{\prime}(E)}\;=\;\frac{\int\exp\left(S_{\rm eff}\right)}{\int O\,\exp\left(S_{\rm eff}\right)}\,.$ (15.3)
Introducing the modified (reweighted) action
$\widetilde{S}\;=\;S_{\rm eff}+\log O\,$, (15.4)
this becomes
$\frac{1}{R(E)}\;=\;\left\langle\frac{1}{O}\right\rangle_{\widetilde{S}}\,$. (15.5)
[…]
Topological stabilization and non-compact excursions
There is a deep geometric reason for this stabilization, naturally understood through the lens of Picard-Lefschetz theory, introduced in QFT by Witten [44], and successfully used in recent QCD simulations [45, 46]. In the complexified phase space, the exact path integral is formally defined on a middle-dimensional integration contour composed of steepest descent manifolds (Lefschetz thimbles). While the deterministic drift of the Complex Langevin Equation naturally targets these thimbles, stochastic noise inevitably kicks the trajectory off these stable manifolds and into adjacent “Stokes wedges,” where the action diverges and the Markov chain suffers runaway excursions.
To counteract this, modern lattice simulations typically rely on computationally expensive algorithmic interventions. As detailed in the review [46], standard approaches involve explicitly deforming the integration contour via holomorphic gradient flow (which incurs a prohibitive $\mathcal{O}(V^{3})$ Jacobian cost), utilizing machine-learning-optimized manifolds, or applying ad-hoc dynamic stabilization and gauge cooling to actively force the system to remain near the thimbles.
Our logarithmic reweighting, however, achieves this stabilization natively for the observable. By shifting the effective action ($\tilde{S}=S_{eff}+\log\mathcal{O}$), we fundamentally alter the topology of the complexified phase space. The associated CLE drift force acquires essential meromorphic poles exactly at the zero-locus of the observable ($\mathcal{O}=0$):
$-\nabla\tilde{S}=-\nabla S_{eff}-\frac{\nabla\mathcal{O}}{\mathcal{O}}$ (15.6)
Crucially, in the Picard-Lefschetz framework, steepest descent flows cannot cross meromorphic poles. By explicitly injecting these poles into the drift dynamics, we puncture the complex manifold. The singular repulsive force acts as an absolute topological obstacle that dynamically corrals the stochastic Markov chain. It safely traps the trajectory within the basin of the physical thimbles and mathematically forbids it from escaping into the runaway regions of the complex plane associated with the zeros of $\mathcal{O}$. This organically enforces dynamic stabilization without the need for computationally devastating Jacobian evaluations.
The intermittent excursions that do occur in our simulation are not numerical instabilities, but rather the exact analytic consequence of the complexified noise exploring the vicinity of this meromorphic pole. As shown in the smooth histogram of the micro-states (see Figure 18), the probability density perfectly matches the analytically predicted $1/|u|^{3}$ power-law fat tails (derived in detail in Appendix D). This exact geometric matching confirms that the finite-sample Cauchy Principal Value is perfectly stable, and that the CLE is genuinely exploring the correct complexified geometry dictated by the topological barrier.
[…]
To prevent the system from exploring these unphysical infinite-volume regimes, we currently impose explicit numerical clipping (hard bounding walls) on the twistor variables. While this pragmatic regularization successfully bounds the metric, there is no a priori mathematical guaranty that the CLE trajectory will not systematically drift outward and continuously collide with these artificial walls, which would artificially truncate the sampled distribution.
Remark 15.1.
It should be noted that the use of explicit numerical clipping (hard bounding walls) to tame the non-compact Liouville directions technically violates the strict holomorphy required for the formal proof of correctness of the Complex Langevin Equation, as it introduces non-analytic boundaries into the Fokker–Planck evolution. A theoretically cleaner approach would involve introducing a globally holomorphic confining potential to the effective action that generates a smooth restoring force. For example, one could augment the action to produce a holomorphic drift force such as $K_{\text{wall}}\propto-\tanh(c(\rho-\rho_{\max}))$. Because the hyperbolic tangent is a meromorphic function of the complex variable $\rho$, this force smoothly pushes the trajectory back toward the physical region $\operatorname{Re}(\rho)<\rho_{\max}$ at large distances without breaking the analyticity of the Langevin evolution. Implementing such holomorphic barriers represents a natural analytical refinement for future high-precision simulations, safely taming the infinite-volume runaway channels while preserving the exact geometric justification of the CLE.
To explicitly defend our simulation against accidental discrete jumps over the topological barriers, we replaced the naive Euler integration with a second-order Stochastic Heun (predictor-corrector) SDE solver. Furthermore, we implemented an adaptive time-step algorithm designed to dynamically brake the CLE evolution upon approaching singular regions. Together, these algorithmic safeguards yield highly stable trajectories that remain predominantly confined to the physical integration cycles (Lefschetz thimbles), consistently producing finite expectation values with rigorously controlled absolute errors.
[…]
Discretization and parallel ensemble
We discretize the boundary circle by a spectral grid (angular nodes) and evaluate the bulk (holographic) Liouville contribution by Gauss–Legendre radial quadrature, leading to a finite-dimensional holomorphic system for the complexified boundary variables evolved by CLE. To reduce sensitivity to non-Gaussian (fat-tailed) fluctuations and to monitor reproducibility, we run an ensemble of independent replicas on the Typhon cluster. In the current production setting, we use 64 nodes (independent seeds) per energy value and pool the results across nodes using robust estimators (median central value and absolute-deviation-based error estimates). The precise implementation and diagnostics are documented in Appendix D.
16 Topological Decomposition of the Meson Spectrum via Information Clustering
16.1 Lefschetz Thimbles and Multiple Spectral Branches
In the planar limit, the exact solution of the Makeenko–Migdal loop equation involves evaluating a path integral over a highly complex configuration space. The effective action landscape of this theory is not dominated by a single, featureless ground state, but rather by a multitude of complex saddle points, or Lefschetz thimbles. Physically, these distinct topological sectors correspond to the multiple branches of the meson spectrum. In our numerical approach, we explore this landscape using stochastic Langevin dynamics. Because each independent simulation is initialized from a completely random configuration, the system naturally falls into the basin of attraction of one of these specific thimbles as it relaxes.
16.2 Topological Trapping and the Failure of Naive Pooling
Due to the presence of immense topological (or simply numerical) action barriers in the complexified space, a given simulation history generally remains confined to its respective thimble for the duration of the run. This dynamical trapping has profound implications for data analysis. If one were to follow standard numerical practice and naively pool all independent stochastic histories at a given energy $E$ to compute a global ensemble average, the procedure would physically mix distinct topological sectors. Consequently, this blind averaging does not yield a single, smooth resolvent curve for $1/\mathcal{R}(E)$; rather, it produces a chaotic, smeared superposition of states. The raw numerical data are actually a collection of distinct histories confined to various thimbles, each tracing out its own unbroken $1/\mathcal{R}(E)$ curve separated by topological barriers.
16.3 Phase-Information as a Topological Fingerprint
To extract the true physical spectrum, we must mathematically untangle these mixed histories and group them by the specific saddle point they occupy. As established in our analysis of the Complex Langevin dynamics in E, the Cartesian observables $\mathrm{Re}[1/\mathcal{O}]$ and $\mathrm{Im}[1/\mathcal{O}]$ suffer from severe infinite-variance Cauchy tails ($1/|u|^{3}$), which generate massive radial excursions. Consequently, attempting to compare the Cartesian coordinates forces one to arbitrarily trim and shift the central moments to prevent the variance from dominating the signal.
We can completely bypass this pathological radial variance by projecting the data onto the complex phase angle $\alpha=\arg(\mathcal{O})$. This maps the entire complexified phase space onto a finite, compact support $\pi,\pi]$, safely absorbing infinite radial fluctuations without discarding a single data point. The probability density function of this phase angle is intrinsically scale-invariant and serves as an exact, bounded topological fingerprint of the underlying Lefschetz thimble. To compare these fingerprints, we compute the symmetric Kullback–Leibler (KL) divergence (the Jensen–Shannon distance) between their finite-support phase distributions. This strict information-theoretic metric allows us to rigorously identify pairs of histories with identical quantum geometries, entirely independent of their macroscopic radial drift.
16.4 Isolating True Mass Poles
This phase-matching provides a precise, unsupervised mechanism for identifying true meson masses. A physical mass corresponds to a strictly downward zero-crossing of the resolvent, where the sequence transitions from $\mathrm{Re}[1/\mathcal{R}(E_{i})]>0$ to $\mathrm{Re}[1/\mathcal{R}(E_{i+1})]<0$. By examining pairs of runs at neighboring energy points that bracket the zero axis and demanding that they possess the absolute minimum symmetric KL distance between their phase distributions, we effectively filter out random topological jumps and numerical noise. This strict requirement of phase-geometry invariance unequivocally identifies the true crossing of the $1/\mathcal{R}(E)$ curve for a particular topological branch.
16.5 Reconstructing the Spectrum via Minimax Clustering
Having isolated a pool of high-confidence zero-crossings, we reconstruct the global spectrum of the different branches. To achieve this, we cluster the crossing events based on a strict minimax distance criterion. The topological distance between any two distinct crossing events—say, a pair $(A,B)$ and a pair (C,D)(C,D)—is defined as the maximum of all four cross-comparisons:
$\Delta=\max\left(d_{JS}(A,C),d_{JS}(A,D),d_{JS}(B,C),d_{JS}(B,D)\right)$ (16.1)
where $d_{JS}$ is the symmetric Kullback–Leibler distance (Jensen–Shannon divergence) between the phase distributions. This mathematically merciless constraint guaranties that a pole is only clustered into a specific sequence if the quantum phase geometry, both before and after the crossing, is practically identical to the rest of the cluster.
16.6 First results of numerical solution of CLE
The CLE was coded into a Mathematica ® package, tested on the laptop and desktop computers, and then run on the Typhon IAS cluster for 48 hours on 64 nodes in parallel, with 24 kernels working in parallel on each node. The code and the results are published on Wolfram cloud, and results are stored on the authors’ Google Drive with open access (see references in the Data section in the end of this paper). Using these Mathematica ® files, one can verify and reproduce our results at a larger scale with higher precision.
Here is the summary of the results (all technical details are described in E).
This unsupervised information-clustering natively reconstructs the spectrum into distinct branches, mapping out different Regge trajectories. Crucially, these trajectories are strictly labeled by an internal quantum number inherent to the Twistor String path integral complex saddle points, serving as a replacement for the “daughter trajectories” of string models.
Remark 16.1.
Note that all the usual quantum numbers are identical for these two branches: this is a scalar $\bar{q}q$ channel, no spin, no flavor quantum numbers specified. In the future, more detailed simulations, spin, parity, and flavor $SU(3)$ representation will be specified, but this internal quantum number will exist on top of other quantum numbers, in the same way as internal excitations of the string (vibrational modes) lead to daughter trajectories after quantization.
Applying this algorithm to our unvarnished stochastic phase data successfully isolated two continuous principal Regge trajectories spanning $14$ and $15$ consecutive negative crossings. The negative crossing of zero between two consecutive points of energy on a grid was identified with accuracy $\Delta E=\pm 1/480\approx\pm 0.002$. Both branches exhibit non-linearity at low energies—reflecting the complex boundary dynamics and mass gaps of the true physical ground state—before transitioning into linear Regge trajectories at higher mass squared. By relying purely on the phase-information geometry, this topological extraction is designed to reveal the true spectrum of Planar QCD.
Figure 15 shows preliminary cluster-scale results for the extracted trajectory $n$ vs. $m_{n}^{2}$. These plots are intended as a proof of concept: the crossings of zero were selected by information clustering of the phase distributions we described above. The precision of higher excitations and the control of systematic effects (step size, resolution, and CLE correctness diagnostics) will be improved in larger-scale runs; the present computation establishes feasibility and provides a concrete computational pipeline for continuum planar-QCD spectroscopy in this framework.
Figure 15: Plot of $n$ vs $m_{n}^{2}$ (in units of the string tension $\sigma$) as a result of Complex Langevin Simulation of quark two point function in the Geometric QCD Theory. These two Regge trajectories were identified by information clustering of phase distributions of $1/O$ for $240$ different energy points between $2.$ and $3.$ on $64$ independent CLE simulations with different initial values. The optimal clustering identified a pair of Regge trajectories in this interval of $m^{2}$
@@ 17.1 Geometric Quantization via Degenerate Lefschetz Thimbles
While the Complex Langevin Equation provides a robust numerical method to evaluate the twistor-string path integral, the holomorphic nature of the effective action reveals a profound mathematical pathway to compute the exact meson spectrum analytically, entirely bypassing stochastic simulation. This approach relies on the asymptotic evaluation of multiple complex integrals via Picard-Lefschetz theory and the singularity theory of differentiable maps.
@@
$\Pi(E)=\int_{\Gamma}\mathcal{D}\Phi_{|z|=1}\,\exp\left(E\,\mathcal{O}(\Phi)-S_{eff}(\Phi)\right)$ (17.1)
where $\Phi_{|z|=1}=\left\{z_{2},\left\{\rho,\xi,\eta,\bar{\xi},\bar{\eta},\gamma,\bar{\gamma}\right\}_{k}\right\}$ encompasses the boundary twistor variables extended analytically into the complexified phase space $\Phi=(\lambda(z),\mu(z))$, and $\Gamma$ is the middle-dimensional integration cycle (the principal Lefschetz thimble) bounding the physical vacuum.
In the semiclassical (WKB) limit, this integral is dominated by the saddle points $\Phi_{\star}(E)$ satisfying the classical boundary equations of motion:
$\nabla S_{eff}(\Phi_{\star})-E\nabla\mathcal{O}(\Phi_{\star})=0$ (17.2)
@@
$\Pi(E)\propto\frac{\exp\left(E\,\mathcal{O}(\Phi_{\star})-S_{eff}(\Phi_{\star})\right)}{\sqrt{\det\mathcal{H}(E)}}$ (17.3)
where the fluctuation operator is $\mathcal{H}(E)=\nabla^{2}S_{eff}(\Phi_{\star})-E\nabla^{2}\mathcal{O}(\Phi_{\star})$.
@@
However, our twistor-string geometry orchestrates a remarkable mathematical conspiracy. Because the integration variables $\Phi$ are coupled complex holomorphic twistors, $\mathcal{H}(E)$ is a complex operator. The vanishing of a single complex eigenvalue intrinsically mandates the simultaneous flattening of exactly two real directions on the Lefschetz thimble (a strict corank-2 degeneracy).
@@
$\det_{\mathbb{C}}\Big[\nabla^{2}S_{eff}(\Phi_{\star})-E_{n}\nabla^{2}\mathcal{O}(\Phi_{\star})\Big]=0$ (17.7)
@@ 17.4 One pole and its symmetries
We may utilize this unbroken space rotation to align the twistor dynamics. Specifically, we can always choose a spatial frame such that the symmetric bilinear combination vanishes:
@@
Because the spatial tangent vectors of the minimal surface are constructed via $f^{\prime}_{k}\propto\lambda^{T}\sigma_{k}\mu$, this strict algebraic constraint forces one component of the tangent vector to be zero everywhere on the worldsheet. This dimensionally reduces the physical minimal surface from a generic 4-dimensional embedding into a 3-dimensional hyperplane, exactly matching the geometric prerequisite for a canonical Helicoid.
@@
The string dynamically centers itself to restore the $U(1)$ symmetry. By locking the pole to the origin, the minimal surface reduces exactly to the canonical, maximally symmetric Helicoid. The remaining computations then proceed in the exact same way as for the canonical Helicoid in Minkowski space. As we derive in a new section in the Appendix, the worldsheet metric following from these canonical twistors $\lambda,\mu$ transforms precisely into the canonical Nambu-Goto rotating string metric. In the text below, we simply use this metric.
17.5 The exact spectrum in the one pole sector
With the twistor saddle point locked to $z_{0}=0$ and the surface constrained to the 3D Helicoid, we repeat the action computation. In this one-pole sector, there are no other terms; the entire worldsheet dynamics are governed by exactly two terms within the Liouville action: the classical area term ($\sigma$) and the kinetic anomaly ($|\nabla\rho|^{2}$).
Because the conformal density is strictly proportional to $1/|z|^{2}$, the kinetic term acts as a rigid, local shift to the effective Helicoid string tension. Evaluating the 1-cycle Minkowski bulk action, we obtain the effective Helicoid energy, complete with the anomalous Lüscher term:
$S^{(1)}_{bulk}=-2\pi^{2}\left(\sigma+\frac{1}{12\pi R^{2}}\right)R^{2}=-2\pi^{2}\sigma R^{2}-\frac{\pi}{6}$ (17.10)
Simultaneously, the boundary term coupling the string to the external energy $E$ integrates to yield the exact phase monodromy per winding:
$\Delta S_{E}=4\pi ER$ (17.11)
Summing these to form the total phase of one physical cycle, we enforce the geometric series resonance condition ($\Delta S_{\text{eff}}=2\pi n$):
$4\pi ER-2\pi^{2}\sigma R^{2}-\frac{\pi}{6}$
$=2\pi n\implies 2ER-\pi\sigma R^{2}-\frac{1}{12}=n$ (17.12)
To locate the physical vacuum, we minimize this effective base action with respect to the spatial radius $R$ to balance the string tension against the injected energy:
$\frac{\partial}{\partial R}\left(2ER-\pi\sigma R^{2}\right)=2E-2\pi\sigma R=0\implies R=\frac{E}{\pi\sigma}$ (17.13)
Substituting this exact saddle-point radius back into the quantization condition, we obtain the exact algebraic spectral equation:
$2E\left(\frac{E}{\pi\sigma}\right)-\pi\sigma\left(\frac{E}{\pi\sigma}\right)^{2}-\frac{1}{12}=n$ (17.14)
Simplifying the squares yields:
$\frac{E^{2}}{\pi\sigma}=n+\frac{1}{12}\implies E^{2}=\pi\sigma\left(n+\frac{1}{12}\right)$ (17.15)
Since we are in the rest frame, the invariant mass squared is $m^{2}=E^{2}$, resulting in the final string spectrum:
$m^{2}=\pi\sigma\left(n+\frac{1}{12}\right)$ (17.16)
We got the famous Lüsher term, without any string models or assumptions. We obtained this term from the conformal anomaly left from the Elfin determinant, which we used to induce MM loop equations. This remarkable coincidence tells us that our solution for planar QCD matches accepted phenomenology as well as the results of the real and numerical experiments. QED! (or rather QCD!).
17.6 QCD as a Geometric Theory
The derivation of the exact spectral equation in the previous section represents a significant shift in our understanding of the large-$N_{c}$ meson spectrum. We have moved from a stochastic description (random surfaces) to a deterministic geometric principle. Let us summarize the key physical insights that have emerged from this analysis.
Our approach avoids the usual “graviton problem” of gravity holography: since our bulk is not dynamical, there are no propagating bulk graviton modes (or other unwanted massless bulk fields) that would have no counterpart in confining QCD.
More precisely, the propagating degrees of freedom in the present formulation are confined to the boundary: (i) the quark loop degrees of freedom, represented by the matrix trace of the Dirac propagator around the boundary loop, and (ii) the holomorphic moduli (twistors) that parametrize the rigid minimal surface. The twistor moduli are specified by their boundary data and extended uniquely into the disk by analytic continuation (Plemelj/Hilbert reconstruction in the spirit of the Penrose transform). Accordingly, bulk quantities such as the Liouville field are not integrated over as independent quantum fields; they are determined functionally by the analytically continued twistor data.
It is also important not to conflate this construction with Witten/Berkovits twistor strings or other topological twistor-string models. In those theories the extension of boundary data into the bulk typically contributes a topological functional (e.g. a WZW term) whose value depends only on the homotopy class of the extension. In contrast, our bulk functional is not topological: its value depends on the specific holomorphic disk selected by the boundary loop, and this non-topological dependence is essential for confinement and for the existence of a mass scale. In this sense, our “rigid twistor string” is best viewed as gauge holography (bulk as a rigid encoding of boundary dynamics), rather than gravity holography or a topological string.
Finally, the observation relating the QCD mass spectrum to the degenerate saddle points of the complexified twistor-string action introduces the concept of localization of the string path integral, but on an entirely different level. Our spectral equation is exactly localized, much like the computation of path integrals in geometric quantization [51], but this is not a familiar fixed point in the moduli space. Instead, the path integral localizes exclusively on degenerate saddle points governed by catastrophe theory. We thus have uncovered a novel, purely geometric mechanism of localization that deterministically dictates the QCD mass spectrum.
17.7 Future Directions: Glueballs, Baryons and Scattering
The formulation of Planar QCD as an Analytic Twistor String opens several avenues for future research:

### v2

@@ 15.5 Balance of quark forces and boundary condition
This result confirms that the linear confinement natively arises directly from the kinematics of the Hodge-dual minimal surface in Minkowski space.
@@ 16.1 Geometric Quantization via Degenerate Lefschetz Thimbles
The holomorphic nature of the effective action reveals a profound mathematical pathway to compute the exact meson spectrum analytically, entirely bypassing stochastic simulation. This approach relies on the asymptotic evaluation of multiple complex integrals via Picard-Lefschetz theory and the singularity theory of differentiable maps.
@@
$\Pi(E)=\int_{\Gamma}\mathcal{D}\mathcal{V}_{|z|=1}\,\exp\left(E\,\mathcal{O}(\mathcal{V})-S_{eff}(\mathcal{V})\right)$ (16.1)
where $\mathcal{V}_{|z|=1}=\left\{z_{2},\left\{\rho,\xi,\eta,\bar{\xi},\bar{\eta},\gamma,\bar{\gamma}\right\}_{k}\right\}$ encompasses the boundary twistor variables extended analytically into the complexified phase space $\mathcal{V}=(\lambda(z),\mu(z))$, and $\Gamma$ is the middle-dimensional integration cycle (the principal Lefschetz thimble) bounding the physical vacuum.
In the semiclassical (WKB) limit, this integral is dominated by the saddle points $\mathcal{V}_{\star}(E)$ satisfying the classical boundary equations of motion:
$\nabla S_{eff}(\mathcal{V}_{\star})-E\nabla\mathcal{O}(\mathcal{V}_{\star})=0$ (16.2)
@@
$\Pi(E)\propto\frac{\exp\left(E\,\mathcal{O}(\mathcal{V}_{\star})-S_{eff}(\mathcal{V}_{\star})\right)}{\sqrt{\det\mathcal{H}(E)}}$ (16.3)
where the fluctuation operator is $\mathcal{H}(E)=\nabla^{2}S_{eff}(\mathcal{V}_{\star})-E\nabla^{2}\mathcal{O}(\mathcal{V}_{\star})$.
@@
However, our twistor-string geometry orchestrates a remarkable mathematical conspiracy. Because the integration variables $\mathcal{V}$ are coupled complex holomorphic twistors, $\mathcal{H}(E)$ is a complex operator. The vanishing of a single complex eigenvalue intrinsically mandates the simultaneous flattening of exactly two real directions on the Lefschetz thimble (a strict corank-2 degeneracy).
@@
$\det_{\mathbb{C}}\Big[\nabla^{2}S_{eff}(\mathcal{V}_{\star})-E_{n}\nabla^{2}\mathcal{O}(\mathcal{V}_{\star})\Big]=0$ (16.7)
@@ 16.3 Poles inside the disk and the monodromies
Remark 16.2.
Localization of the spectrum vs. localization of path integral
It is crucial to distinguish between the exact spectral condition and the exact computation of the path integral. In the complexified functional space, there are infinitely many directions away from the critical saddle point (where a complex eigenvalue vanishes). In the framework of Stochastic Quantization and the Complex Langevin Equation (CLE), the stochastic trajectory is effectively exploring a Lefschetz Thimble—the stable manifold of the gradient flow [5]. The trajectory passes through the saddle and proceeds along an infinitely long, "flat" minimal valley, while stochastic noise causes it to fluctuate against the steep potential walls surrounding this valley.These transverse fluctuations contribute to the pre-exponential factor (comprising the functional determinant and higher-order expansion terms). However, as established in the topological study of Landau singularities by [41], the global properties of the integral are determined by the monodromy of the integration cycle. The trajectory must wind infinitely many times around the logarithmic branching point of the effective action $S_{eff}(z)$.Each circuit around this singularity shifts the Minkowski action by a fixed amount, $\mathrm{i}\Delta S_{eff}$. The condition for a pole $E=E_{n}$ in the total amplitude corresponds to the constructive interference of these windings, leading to the quantization condition:
$\Delta S_{eff}=2\pi n$ (16.8)
This monodromy is a topological invariant of the flow and is independent of the "shape" of the valley sides or the specific stochastic time, which are non-universal. This separation of local geometry from global topology is a manifestation of the spectral flow and the index theorems of [2], where the "integer" part of the quantization (the pole) remains robust against the local deformations of the path caused by Langevin noise. As demonstrated in our WKB analysis of the helicoid, this mechanism ensures that the energy spectrum is localized by the classical monodromy, even while the path integral remains delocalized by the CLE.
@@ 16.4 One pole and its symmetries
In a conventional approach modeling a generic “rubber band” string, one might utilize this unbroken space rotation to align the twistor dynamics such that the symmetric bilinear combination vanishes:
@@
Because the spatial tangent vectors of the minimal surface are constructed via $f^{\prime}_{k}\propto\lambda^{T}\sigma_{k}\mu$, this strict algebraic constraint would force one component of the tangent vector to be zero everywhere on the worldsheet. This would dimensionally reduce the physical minimal surface from a generic 4-dimensional embedding into a 3-dimensional hyperplane.
However, here is where the profound difference between the conventional “rubber band” string and our twistor formulation manifests itself. Instead of crushing the surface into 3D, we enforce the exact chiral (Hodge-dual) configuration. Specifically, we freeze one of the chiral sectors by setting the right-handed twistor to a constant basis spinor, such as $\mu=(1,0)^{T}$. This choice explicitly breaks the 3D embedding constraint, causing the minimal surface to genuinely span a 4-dimensional space and flawlessly evade the finite-area self-intersection theorems. Because macroscopic QCD is strictly parity-conserving, the physical target-space configuration cannot remain purely left- or right-chiral. The true physical minimal surface is given by our exact prescription: the parity-even arithmetic mean of the left-chiral (Self-Dual or SD) and right-chiral (Anti-Self-Dual or ASD) surfaces. As derived in Appendix F, this symmetrization rigorously projects out the imaginary geometric artifacts of the 4D embedding, yielding a strictly real Nambu-Goto string geometry whose fundamental classical area is exactly halved relative to the conventional generic string.
@@
The string dynamically centers itself to restore the $U(1)$ symmetry. By locking the pole to the origin, the minimal surface reduces exactly to the maximally symmetric canonical Helicoid. The remaining computations then proceed utilizing the arithmetic mean of the SD and ASD areas. In the text below, we simply use this symmetrized Hodge-dual metric.
16.5 Quark Spin Monodromy and the Unified Regge Spectrum
The exact meson spectrum is completed by accounting for the discrete spin monodromy of the boundary quarks. As algebraically demonstrated in F, the discretization of the quark loop into $2L$ steps yields a path-ordered Dirac trace $T\propto(-1)^{L}$. This phase factor provides an intrinsic topological index that modifies the constructive interference condition in the worldsheet amplitude, selecting specific sectors of the geometric series $\sum\exp\left(iS_{k}\right)$.
The resulting Bohr-Sommerfeld resonance condition is $\Delta S_{\text{eff}}=\pi n_{BS}$. Based on the trace $T$, physical light mesons ($\pi$ and $\rho$) are found to occupy the even-$n_{BS}$ sectors. Defining the worldsheet topological quantum number $n=n_{BS}$ and evaluating the 1-cycle Minkowski bulk action for the symmetrized chiral Helicoid (Appendix F), we obtain the unified mass-squared formula:
$m^{2}=\frac{\pi\sigma_{e,o}}{4}\left(n+\frac{1}{12}\right)$ (16.11)
For even $L$, the projector supports quantization with $n\equiv 0\pmod{4}$ and $n\equiv 2\pmod{4}$. To map these topological sectors to the macroscopic meson spectrum and define the intercepts $\alpha(0)=J|_{m^{2}=0}$, we adopt the semi-classical WKB mapping $n=4J+q$: